OpenStudy (anonymous):

I have a mole to mole conversion for chemistry that I do not understand how to work. You have a sample of dinitrogen pentoxide that contains 40 g of oxygen. What is the mass of your sample of dinitrogen pentoxide?

OpenStudy (rushwr):

SO that iss N2O5

OpenStudy (rushwr):

N2O5 to oxygen ratio is 1:5 which means if we have 5 moles of oxygen here we will find 1 mole N2O5 in this compound .

OpenStudy (anonymous):

\[40 g O \ \frac{ 1 mol O }{ 16 g O }\ \frac{ 1 mol N2O5 }{ 5 mol O } \frac{ 108 g N2O5 }{ 1 mol N2O5 } = 54 g N2O5\]

OpenStudy (rushwr):

SO first u have to find the no. of moles of oxygen moles = mass divided by molar mass moles of O2 = 40/16 = 2.5 moles Earlier we said N2O5 to oxygen ratio = 1:5 We have 2.5 moles of Oxygen , then to find the no. of moles of N2O5= 2.5/5 = 0.5 moles Using the previous equation we can take mass= molar mass*moles mass of N2O5 = 108 * 0.5 = 54g of N2O5